# HyResPINNs: Adaptive Hybrid Residual Networks for Learning Optimal Combinations of Neural and RBF Components for Physics-Informed Modeling

### v1

@@ 4.2 Darcy Flow
4.2.1 Annulus Domain with Smooth Coefficients
We initially showcase the convergence rate of our method by employing smooth manufactured solutions. We consider two scenarios: one within an annular 2D domain, and another within a 3D domain formed by extruding the 2D geometry in the z-direction, resulting in a cylinder with a height of two. We present a convergence study analyzing the error rate as the number of points increases. The exact solutions are set as,
$u(x,y)=\sin(x)\sin(y),\;\;\;\;u(x,y,z)=\sin(x)\sin(y)\sin(z),$ (16)
for the 2D and 3D cases, respectively. By substituting these exact solutions into the model problem we define the source term and the boundary data. We solve the discrete problem with either Dirichlet or Neumann boundary conditions. Convergence results are shown in Figure 7 for the two and three dimensional problems and demonstrate that HyResPINNs  remain robust to boundary condition type, size of training set (achieves lowest errors compared to the baselines for fewer training points), and problem dimension.
4.2.2 2D Box Domain with Rough Coefficients
A defining feature of methods for solving the Darcy Flow equation 15, lies in their proficiency in accurately depicting the flux variable in scenarios where the coefficients $\mu$ are discontinuous. In these instances, the flux’s normal component maintains continuity across material interfaces, whereas the tangential component may exhibit discontinuities. In contrast, collocated methods like Galerkin, stabilized Galerkin, and least-squares finite elements often fail to replicate this physical behavior, typically resulting in oscillations at the interface. In this section, we demonstrate that HyResPINNs  successfully provides physically accurate flux approximations for problems characterized by discontinuous coefficients.
The initial example presented is the well-documented "five strip problem" Nakshatrala et al. (2006); Masud & Hughes (2002), which serves as a conventional manufactured solution test to evaluate a method’s capability to preserve the continuity of normal flux. The prescribed exact solution on domain $\Omega=[0,1]^{2}$ is given by,
@@ 4.2.1 Annulus Domain with Smooth Coefficients
with different $\mu_{i}$ on each $\Omega_{i}$ such that $\mu_{1}=16,\;\;\mu_{2}=6,\;\;\mu_{3}=1,\;\;\mu_{4}=10,\;\;\mu_{5}=2$. We solve the discrete problem with Neumann boundary conditions. Convergence results are shown in Figure 8 and show that HyResPINNs  provide accurate solutions, albeit with larger training sizes.
5 Conclusion
In this work, we introduced HyResPINNs, a novel class of physics-informed neural networks that incorporate adaptive hybrid residual blocks combining the strengths of standard neural networks and radial basis function (RBF) networks. Our architecture effectively captures continuous and discontinuous features by leveraging smooth and non-smooth function approximators. The adaptive combination parameters within each block allow the model to balance the contributions of neural and RBF components during training. Furthermore, including Wendland kernels enhances the model’s ability to handle sharp transitions while maintaining computational efficiency. Our experiments demonstrate that HyResPINNs  outperform traditional PINNs and state-of-the-art methods in accuracy and robustness, particularly for problems involving mixed smooth and non-smooth regions. This work increases the flexibility and generalizability of PINNs by bridging the gap between classical numerical methods and DNN-based approaches for solving PDEs.
Appendix A Experimental Design
We use mini-batch gradient descent, where collocation points are randomly sampled during training iteration, for all Allen-Cahn experiments, and we use full-batch gradient descent (but varying training set size) for all other experiments. We use the Adam optimizer Kingma & Ba (2015), and follow the learning rate schedule of Wang et al. (2023) which starts with a linear warm-up phase of $5,000$ iterations, starting from zero and gradually increasing to $10^{-3}$, followed by an exponential decay at a rate of $0.9$. Following the best practices described in Wang et al. (2023), we also employ a learning rate annealing algorithm Wang et al. (2021a; 2023) to balance losses and causal training Wang et al. (2022a; 2023) to mitigate causality violation in solving time-dependent PDEs and apply exact periodic boundary conditions Dong & Ni (2021) when applicable. We use the hyperbolic tangent activation functions and initialize each network’s parameters using the Glorot normal scheme, unless otherwise specified. We ran five random trials for each test, and report the mean values achieved in each plot and table.

### v2

@@ B.2.1 Smooth Coefficients
B.2.2 Rough Coefficients
Figure 10: (Left) True solution on the 2D box domain. (Middle) True transverse flux. (Right) True longitudinal flux.
The prescribed exact solution on domain $\Omega=[0,1]^{2}$ is given by,
@@ B.2.2 Rough Coefficients
with different $\mu_{i}$ on each $\Omega_{i}$ such that $\mu_{1}=16,\;\;\mu_{2}=6,\;\;\mu_{3}=1,\;\;\mu_{4}=10,\;\;\mu_{5}=2$.
B.3 Kuramoto–Sivashinsky Equation
We use the 1D Kuramoto–Sivashinsky equation of the form:
$\frac{\partial u}{\partial t}+\alpha u\frac{\partial u}{\partial x}+\beta\frac{\partial^{2}u}{\partial x^{2}}+\gamma\frac{\partial^{4}u}{\partial x^{4}}=0,$ (19)
subject to periodic boundary conditions and an initial condition
$u(0,x)=u_{0}(x).$ (20)
The parameters $\alpha$, $\beta$, and $\gamma$ control the dynamical behavior of the equation. We use the configurations: $\alpha=100/16$, $\beta=100/162$, and $\gamma=100/164$ for chaotic behaviors.
Appendix C Experimental Design
We use mini-batch gradient descent, where collocation points are randomly sampled during training iteration, for all Allen-Cahn experiments, and we use full-batch gradient descent (but varying training set size) for all other experiments. Each full-batch collocation point sets were generated using the efficient Poisson sampling technique from Shankar (2017). We use the Adam optimizer Kingma & Ba (2015), and follow the learning rate schedule of Wang et al. (2023) which starts with a linear warm-up phase of $5,000$ iterations, starting from zero and gradually increasing to $10^{-3}$, followed by an exponential decay at a rate of $0.9$. Following the best practices described in Wang et al. (2023), we also employ a learning rate annealing algorithm Wang et al. (2023) to balance losses and causal training Wang et al. (2022a; 2023) to mitigate causality violation in solving time-dependent PDEs and apply exact periodic boundary conditions Dong & Ni (2021) when applicable. We use the hyperbolic tangent activation functions and initialize each network’s parameters using the Glorot normal scheme, unless otherwise specified. We ran five random trials for each test, and report the mean values achieved in each plot and table.